i need help soving a substitution problom
2x-3y=-14
3x-y=-7

Answer by Alan3354(69443)

2x-3y=-14
3x-y=-7
y = 3x+7 (from the 2nd eqn)
Sub into the 1st eqn
----------
2x-3y=-14
2x - 3(3x+7) = -14
2x - 9x - 21 = -14
-7x = 7
x = -1
-----------
3x-y=-7
3*(-1) - y = -7
-3 - y = -7
y = 4